Two refills for a ball point pen are selected at random for a box that contains 55

blue, 3 red and 3 green refills. If X is the number of blue refills and Y is the number

of red refills selected .Find the joint probability function f (x, y) and p[(x,y)"\isin"A ], where A is the region {(x,y): x+y<=1}


1
Expert's answer
2021-08-24T19:13:06-0400

The total number of refills is: "55+3+3=61". The probability to get a blue refill is: "p_1=\frac{55}{61}". For red refills we have: "p_2=\frac{3}{61}". The joint probability function (probability mass function) "f(x,y)" is given by: "f(x,y)=p_{X,Y}(x,y)=P(X=x\wedge Y=y)". Suppose that we selected "2" refills. There are "m=C_{61}^2=\frac{61!}{2!(61-2)!}=\frac{60\cdot61}{2}=30\cdot61=1830" ways to do this. Suppose that we selected "2" refills and among them there are "x\leq2" blue refills and "y\leq2" red refills. In addition, since 2 refills are selected. "x+y\leq2" It is possible to do in "n=C_{55}^xC_{3}^yC_{3}^{2-x-y}=\frac{55!}{x!(55-x)!}\frac{3!}{y!(3-y)!}\frac{3!}{(2-x-y)!(1+x+y)!}" ways. Thus, we get: "f(x,y)=p_{X,Y}(x,y)=P(X=x\wedge Y=y)=\frac{n}{m}=\frac{C_{55}^xC_{3}^yC_{3}^{2-x-y}}{C_{61}^2}=\frac{\frac{55!}{x!(55-x)!}\frac{3!}{y!(3-y)!}\frac{3!}{(2-x-y)!(1+x+y)!}}{1830}".

The probability "p[(x,y)\in A]", where "A=\{(x,y):x+y\leq1\}" is: "p[(x,y)\in A]=p(X=1,Y=0)+p(X=0,Y=1)=\frac{C_{55}^1C_{3}^0C_{3}^{1}}{C_{61}^2}+\frac{C_{55}^0C_{3}^1C_{3}^{1}}{C_{61}^2}=\frac{55\cdot3}{1830}+\frac{3\cdot3}{1830}=\frac{174}{1830}\approx0.095"